A box contains 12 white and 8 black marbles. Two balls are drawn out randomly from
the box without replacement. Let X denote the number of white balls drawn out.
a. Construct the probability distribution of X.
b. Find mean and variance of X using the following formula
? = E (X) = ∑ ? . ?(?) ?
?(?2) = ∑ ?2
. ?(?) ?
?2 = ???(?) = ?(?2) − (?)2
Answer #1

a.~We~can~draw~2~balls~from~20~balls~in~\binom{20}{2}~ways.~then\\\\ The~probability~distribution~of~X~is\\\\ P(X=0)=\frac{\binom{8}{2}}{\binom{20}{2}}=0.14737\\\\ P(X=1)=\frac{\binom{8}{1}\binom{12}{1}}{\binom{20}{2}}=0.50526\\\\ P(X=2)=\frac{\binom{12}{2}}{\binom{20}{2}}=0.34737\\\\ P(X=x)=0~otherwise\\\\ b.~\mu=E(X)=0*0.14737+1*0.50526+2*0.34737=1.2\\\\ E(X^2)=0^2*0.14737+1^2*0.50526+2^2*0.34737=1.89474\\\\ \sigma^2=Var(X)=E(X^2)-E^2(X)=1.89474-1.2^2=0.45474